I have the following problem; rewrite this equation as a funtion of x; the eqaution is -3x + 4 = 11. If I'm doing this right I get as far as y = -3x + 11 and then I get competely lost, could you please explain the next step to me. I know I'm forgetting an operation, but can't remember what. Thank You.
Answer by checkley71(8403)

-3X+4=11 OR subtracting 4 from both sides we get -3X+4-4=11-4 OR -3x=11-4 OR
-3X=7 OR dividing both sides by -3 we get -3X/-3=7/-3 OR X=7/-3 OR x=-7/3
PROOF -3(-7/3)+4=11 OR 7+4=11 OR 11=11